For interval x € [-1.0) , f(x)=x-1 & for interval x€ [0,1] , f(x)= .x^2 g(x) = sin x h(x) = f(|g(x)|) + |f(g(x))| Using the above data, show that h(x) is discontinuous at 0 in the interval [-1,1] .




For  interval  x €  [-1.0) , f(x)=x-1  &   for  interval  x€ [0,1] , f(x)= .x^2


g(x) =  sin x


h(x) = f(|g(x)|) + |f(g(x))|


Using   the  above  data, show that h(x) is  discontinuous  at 0  in the interval  [-1,1]  .




Grade:

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Vivek,

 

Remember -                               | x,  x≥0
                                  |x| =      
                                               | -x,  x<0       and   [ mod.(x-y) ≥ mod. x - mod. y ]

For x€ [0,1] , f(x)= .x^2 so,  RHL : h(x) =  (mod. sinx )2 + mod. sin2x

                                                                     =   sin2x + sin2x

Lim h -0 we have h(x) = 0


for  x €  [-1.0) , f(x)=x-1 LHL : h(x) = mod. sinx -1 + mod. ( sinx -1 )

                                                                  =  - sinx -1 +mod. sinx - mod. 1

                                                                  =    - sinx -1 - sinx -1

lim h - 0 , we have h(x) = -2

 

As, LHL not equal to RHL so discontinous at 0

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay